# STI-Bench: Are MLLMs Ready for Precise Spatial-Temporal World Understanding?

### v1

@@ Abstract
The use of Multimodal Large Language Models (MLLMs) as an end-to-end solution for Embodied AI and Autonomous Driving has become a prevailing trend. While MLLMs have been extensively studied for visual semantic understanding tasks, their ability to perform precise and quantitative spatial-temporal understanding in real-world applications remains largely unexamined, leading to uncertain prospects. To evaluate models’ Spatial-Temporal Intelligence, we introduce STI-Bench, a benchmark designed to evaluate MLLMs’ spatial-temporal understanding through challenging tasks such as estimating and predicting the appearance, pose, displacement, and motion of objects. Our benchmark encompasses a wide range of robot and vehicle operations across desktop, indoor, and outdoor scenarios. The extensive experiments reveals that the state-of-the-art MLLMs still struggle in real-world spatial-temporal understanding, especially in tasks requiring precise distance estimation and motion analysis.
@@ 3.2 Benchmark Construction
Dim. Meas. Disp. & P.L. Speed & Acc. Spatial Relation Ego Orient. Traj. Desc. 3D Video Grounding Pose Est.
@@
GPT-4o[30] 6 28.3 20.1 20.6 26.9 48.0 22.2 35.9 21.8 42.8
Gemini-2.0-Flash[35] 4 35.6 30.8 25.1 32.6 48.0 20.5 48.7 28.1 59.1
Gemini-2.5-Pro[35] 2 40.5 32.9 30.8 34.0 53.4 55.7 42.9 36.2 52.5
Claude-3.7-Sonnet[2] 3 37.2 32.5 25.4 34.4 48.3 26.5 42.3 34.7 57.2
@@
Qwen2.5-VL-72B[5] 1 40.8 34.6 28.1 36.1 45.5 48.0 46.2 35.8 60.5
InternVL2.5-78B[13] 5 29.3 26.0 20.7 27.0 52.5 19.7 28.4 25.6 45.1
VideoLLaMA3-7B[45] 7 27.4 26.3 24.3 24.5 39.7 21.1 24.4 27.8 32.5
@@
Claude-3.7-Sonnet 39.27 35.84 35.94 37.17
Gemini-2.0-Flash 39.97 34.22 30.07 35.61
Gemini-2.5-Pro 48.02 35.36 36.92 40.52
GPT-4o 36.95 21.27 26.16 28.25
InternVL2.5 33.39 28.85 23.79 29.34
Qwen2.5-VL 48.23 35.48 36.70 40.84
VideoLLaMA3-7B 26.99 28.79 25.12 27.37
@@ 4.1 Settings
We conduct a thorough evaluation of leading MLLMs from diverse model families, focusing on both proprietary and open-source solutions. Specifically, we assess the performance of four proprietary models, GPT-4o[30], Gemini-2.0-Flash[35], Gemini-2.5-Pro[35], and Claude-3.7-Sonnet[2], as well as several representative open-source MLLMs that have undergone specialized video-related training, including Qwen2.5-VL-72B[5], InternVL2.5-78B[13] and VideoLLaMA3-7B[45].
To ensure a consistent temporal sampling strategy across videos, we sample frames at 1 fps for both input and output. Our benchmark tasks are presented in a multiple-choice format with five possible answers, hence a random guess baseline yields a 20% accuracy. We measure each model’s accuracy by directly comparing the model’s selected answer with the ground truth, without employing any additional external models or annotations for performance evaluation.
@@ 4.2 Main Results
As shown in Table 2, we present a comprehensive evaluation of various MLLMs on STI-Bench. Overall, Qwen2.5-VL-72B achieves the highest accuracy of 40.8% among all tested models, slightly outperforming Gemini-2.5-Pro (40.5%). While these results significantly exceed the random guess baseline (20%), they still highlight substantial room for improvement in spatial-temporal understanding.
When analyzing performance across different scene types, we observe consistent patterns. All models perform better in outdoor scenarios (Qwen2.5-VL: 48.2%, Gemini-2.5-Pro: 48.0%) compared to indoor (35.5%, 35.4%) and desktop environments (36.7%, 36.9%). This suggests that models might have been exposed to more outdoor video content during training, or that outdoor scenes often provide clearer visual cues for spatial relationships.
Task-specific performance reveals particularly challenging areas. Most models struggle significantly with Displacement & Path Length estimation (best: 31.6% by Qwen2.5-VL) and Dimensional Measurement (best: 34.6% by Qwen2.5-VL), both of which require precise quantitative spatial understanding. In contrast, models demonstrate stronger capabilities in Pose Estimation (best: 60.5% by Qwen2.5-VL) and Spatial Relation tasks (best: 55.4% by Qwen2.5-VL).
Notably, open-source models like Qwen2.5-VL and InternVL2.5 demonstrate competitive performance compared to proprietary models, with Qwen2.5-VL even slightly outperforming Gemini-2.5-Pro on several metrics. However, smaller models like VideoLLaMA3-7B (27.4% overall) still lag significantly behind their larger counterparts.
It is important to emphasize that even the best-performing models achieve only about 40% accuracy on our benchmark, which, while twice the random guess baseline, remains far from the reliability required for real-world embodied AI or autonomous driving applications. These results indicate that current MLLMs, despite their impressive capabilities in general visual understanding, still struggle with precise spatial-temporal intelligence tasks central to embodied applications.
Figure 6: The simplified thought process examples of Gemini 2.5 Pro.
@@ 4.3 Experimental Analysis
Given that Gemini-2.5-Pro is a multi-modal reasoning model with detailed thinking processes and ranks second-best among all tested models (best among proprietary models), we select it as a representative for in-depth analysis. The simplified thought process examples is presented in Figure 6.
Our analysis of Gemini-2.5-Pro reveals several key characteristics of current state-of-the-art MLLMs’ spatial-temporal understanding capabilities. Overall, the model achieves 40.52% accuracy. Performance is notably stronger in outdoor scenarios (48.02%) compared to desktop (36.92%) and indoor environments (35.36%). This performance disparity suggests that the model’s training data likely emphasized outdoor scenes and larger-scale understanding, or that outdoor environments typically provide clearer visual cues for spatial reasoning.
When examining task-specific performance, we observe that Gemini-2.5-Pro demonstrates stronger capabilities in orientation and spatial relationship tasks. It achieves the highest accuracy in Ego-Centric Orientation (55.74%) and Spatial Relation tasks (53.42%), followed by Pose Estimation (52.53%). However, the model struggles significantly with tasks requiring precise quantitative estimation, particularly Displacement & Path Length (30.81%), Dimensional Measurement (32.87%), and Speed & Acceleration (34.04%).
@@ 5 Conclusion
We introduced STI-Bench, a comprehensive benchmark to assess MLLMs’ spatial-temporal understanding through over 300 real-world videos and 2,000 QA pairs of robot desktop, indoor, and outdoor scenarios, which reveals significant limitations in current MLLMs’ spatial-temporal understanding capabilities, with even top-performing models achieving only 40-48% accuracy. Models particularly struggle with precise quantitative tasks like dimensional measurement. Our analysis identifies three key weaknesses: inaccurate spatial quantification, flawed temporal dynamics understanding, and weak cross-modal integration. These findings emphasize the substantial gap between current capabilities and the reliability needed for embodied AI and autonomous driving applications. STI-Bench provides a valuable framework for evaluating and improving MLLMs’ ability to understand the physical world—essential for developing the next generation of embodied intelligent systems.

### v2

@@ Abstract
The use of Multimodal Large Language Models (MLLMs) as an end-to-end solution for Embodied AI and Autonomous Driving has become a prevailing trend. While MLLMs have been extensively studied for visual semantic understanding tasks, their ability to perform precise and quantitative spatial-temporal understanding in real-world applications remains largely unexamined, leading to uncertain prospects. To evaluate models’ Spatial-Temporal Intelligence, we introduce STI-Bench, a benchmark designed to evaluate MLLMs’ spatial-temporal understanding through challenging tasks such as estimating and predicting the appearance, pose, displacement, and motion of objects. Our benchmark encompasses a wide range of robot and vehicle operations across desktop, indoor, and outdoor scenarios. The extensive experiments reveal that the state-of-the-art MLLMs still struggle in real-world spatial-temporal understanding, especially in tasks requiring precise distance estimation and motion analysis. Paper Page: https://mira-sjtu.github.io/STI-Bench.io/
@@ 3.2 Benchmark Construction
Dim. Meas. Spatial Relation 3D Video Grounding Disp. & P.L. Speed & Acc. Ego Orient. Traj. Desc. Pose Est.
@@
GPT-4o[31] 5 34.8 24.9 49.6 28.1 27.6 36.0 30.3 36.8 51.3
Claude-3.7-Sonnet[2] 3 39.4 31.8 49.0 36.3 29.0 36.9 27.0 41.0 62.7
Gemini-2.0-Flash[36] 4 38.7 33.7 50.0 33.7 32.7 34.4 15.1 48.7 62.4
Gemini-2.5-Pro[36] 2 40.9 34.2 53.4 32.3 32.4 34.3 44.9 52.0 58.4
@@
VideoLLaMA3-7B[47] 8 26.9 25.6 36.3 24.3 22.4 27.5 23.8 28.2 31.7
MiniCPM-V-2.6[45] 8 26.9 27.7 44.5 29.0 19.0 25.7 7.0 30.8 35.6
InternVL2.5-78B[13] 7 28.4 22.1 38.9 26.0 23.2 30.2 26.5 32.4 36.1
VideoChat-Flash[26] 6 34.1 27.9 21.2 29.7 27.4 22.3 64.7 45.8 46.4
Qwen2.5-VL-72B[5] 1 41.3 32.2 46.2 39.6 28.3 35.6 50.3 44.9 61.0
@@
VideoLLaMA3-7B 26.61 27.63 25.86 26.89
MiniCPM-V-2.6 27.87 26.13 26.60 26.89
InternVL2.5 34.11 25.89 23.81 28.39
VideoChat-Flash 36.12 35.32 27.53 34.08
GPT-4o 39.12 34.45 27.41 34.82
Claude-3.7-Sonnet 43.43 36.53 37.41 39.35
Gemini-2.0-Flash 40.71 37.57 37.41 38.74
Gemini-2.5-Pro 47.48 38.55 33.09 40.88
Qwen2.5-VL 49.24 35.63 36.23 41.32
@@ 4.1 Settings
We conduct a thorough evaluation of leading MLLMs from diverse model families, focusing on both proprietary and open-source solutions. Specifically, we assess the performance of four proprietary models, GPT-4o[31], Gemini-2.0-Flash[36], Gemini-2.5-Pro[36], and Claude-3.7-Sonnet[2], as well as several representative open-source MLLMs that have undergone specialized video-related training, including Qwen2.5-VL-72B[5], InternVL2.5-78B[13], MiniCPM-V-2.6[45], VideoChat-Flash[26] and VideoLLaMA3-7B[47].
Considering the stability of open-source models and the API limitations of proprietary models, we uniformly sample 30 frames from the video for each record and explicitly indicate the sampling FPS (Frames Per Second) for the current video within the prompt. An exception is made for Claude-3.7-Sonnet, for which only 20 frames are sampled due to its API constraints. Our benchmark tasks are presented in a multiple-choice format with five possible answers, hence a random guess baseline yields a 20% accuracy. We measure each model’s accuracy by directly comparing the model’s selected answer with the ground truth, without employing any additional external models or annotations for performance evaluation.
Figure 6: The simplified thought process examples of Gemini 2.5 Pro.
@@ 4.2 Main Results
As shown in Table 2, we present a comprehensive evaluation of various MLLMs on on STI-Bench. Overall, Qwen2.5-VL-72B achieves the highest average accuracy of 41.3% among all tested models, slightly outperforming Gemini-2.5-Pro (40.9%). While these results significantly exceed the random guess baseline (20% ), they still highlight substantial room for improvement in spatial-temporal understanding.
When analyzing performance across different scene types, we observe varied patterns. Qwen2.5-VL demonstrates significantly better performance in outdoor scenarios (49.2%) compared to its indoor (35.6%) and desktop performance (36.2%). Conversely, Gemini-2.5-Pro shows its strongest performance in indoor environments (38.6%), followed by outdoor (47.5%) and then desktop scenes (33.1%). Other models like Claude-3.7-Sonnet and Gemini-2.0-Flash show strength in desktop scenarios (both 37.4%). This suggests varying model specializations, possibly influenced by training data distribution or architectural differences in handling different environmental cues.
Task-specific performance reveals particularly challenging areas. Models generally struggle with tasks requiring precise quantitative estimation. For instance, the best performance on Displacement & Path Length is achieved by Gemini-2.0-Flash (32.7%), and on Dimensional Measurement by Gemini-2.5-Pro (34.2%). Speed & Acceleration also presents difficulties, with Claude-3.7-Sonnet leading at 36.9%. In contrast, models demonstrate stronger capabilities in Pose Estimation (best: 62.7% by Claude-3.7-Sonnet) and Spatial Relation tasks (best: 53.4% by Gemini-2.5-Pro). Notably, VideoChat-Flash achieves the highest score in Ego Orientation (64.7%) despite its lower overall average.
Among open-source models, Qwen2.5-VL stands out, demonstrating highly competitive performance and achieving the top overall rank. However, other open-source models like InternVL2.5, VideoLLaMA3-7B, and MiniCPM-V-2.6 still lag significantly behind the top performers and larger counterparts.
It is important to emphasize that even the best-performing model, Qwen2.5-VL, achieves only 41.3% average accuracy on our benchmark. While this is considerably better than random guessing, it remains far from the reliability required for demanding real-world applications like embodied AI or autonomous driving. These results indicate that current MLLMs, despite their impressive general visual understanding capabilities, still require significant advancements in precise spatial-temporal intelligence for embodied tasks.
@@ 4.3 Experimental Analysis
Given that Gemini-2.5-Pro exhibits strong multi-modal reasoning capabilities, provides detailed thinking processes (as exemplified in Figure 6), and ranks second-best overall (best among proprietary models), we select it as a representative for in-depth analysis.
Our analysis of Gemini-2.5-Pro reveals several key characteristics of current state-of-the-art MLLMs’ spatial-temporal understanding capabilities. Overall, the model achieves 40.9% average accuracy. Performance varies across scene types, with the strongest results in indoor scenarios (38.6%), followed by outdoor (47.5%) and desktop environments (33.1%). This performance disparity suggests that the model’s training data likely emphasized outdoor scenes and larger-scale understanding, or that outdoor environments typically provide clearer visual cues for spatial reasoning.
When examining task-specific performance for Gemini-2.5-Pro (Table 2), we observe its strongest capabilities in tasks involving relative understanding and object state. It achieves its highest score in Pose Estimation (58.4%), followed by Spatial Relation (53.4%, also the best overall score for this task among all models). Its performance in Ego Orientation is 44.9%. Interestingly, while tasks requiring precise quantitative estimation generally show lower absolute scores across models, Gemini-2.5-Pro performs very competitively or even leads in these areas: it achieves the top score in Dimensional Measurement (34.2%) and the second-best score in Displacement & Path Length (32.4%). Its score for Speed & Acceleration is 34.3%. This indicates that while the absolute accuracy in precise estimation tasks remains a challenge, Gemini-2.5-Pro shows comparatively strong capabilities in these domains relative to other models.
@@ 5 Conclusion
We introduced STI-Bench, a comprehensive benchmark to assess MLLMs’ spatial-temporal understanding through over 300 real-world videos and 2,000 QA pairs of robot desktop, indoor, and outdoor scenarios, which reveals significant limitations in current MLLMs’ spatial-temporal understanding capabilities, with even the top-performing models achieving no more than 50% accuracy. Models particularly struggle with precise quantitative tasks like dimensional measurement. Our analysis identifies three key weaknesses: inaccurate spatial quantification, flawed temporal dynamics understanding, and weak cross-modal integration. These findings emphasize the substantial gap between current capabilities and the reliability needed for embodied AI and autonomous driving applications. STI-Bench provides a valuable framework for evaluating and improving MLLMs’ ability to understand the physical world—essential for developing the next generation of embodied intelligent systems.
Acknowledgements
We thank the data annotation team at MolarData for their support, as well as the volunteer contributors Zixin Liu, Junjie Ruan, Lingxi Xia, Zixuan Huang, Tianrui Wan, Qingcheng Wei, Yujie Yao, Shangyang Dong, Zhuofan Zeng, and Yiming Li for their valuable work as human annotators and evaluators.
Appendix A Technical Details for Benchmark Construction and Analysis
A.1 Dataset Interface
When searching for datasets, we found that apart from RGB frame data, other data formats are inconsistent. For example, we noticed that it does not provide 3D bounding box information, so we could only use it to compute camera displacement, velocity, and other related physical quantities. In contrast, Waymo includes such information, which is why we aim to align different datasets into a unified format. Therefore, we unify different types of data by converting them into frame-by-frame instance-level 3D point clouds or bounding box annotations, along with per-frame camera parameters, enabling us to obtain all the required information and process it in a consistent manner.
A.2 Ground Truth Computation
In STI-Bench, the ground truth annotations are derived from multi-modal data, including frame-by-frame camera intrinsics/extrinsics and precise 3D point cloud annotations. We utilize three publicly available datasets (Omni6DPose, ScanNet, Waymo) to cover different scenarios (desktop, indoor, outdoor). Below, we provide eight benchmark tasks with unified notations and formulas.
A.2.1 Static Understanding
Dimensional Measurement
Let $l_{x}$, $l_{y}$, $l_{z}$ denote the dimensions (length, width, height) of an object along the $x$, $y$, and $z$ axes:
$l_{x}$ $=x_{\max}-x_{\min},$ (1)
$l_{y}$ $=y_{\max}-y_{\min},$
$l_{z}$ $=z_{\max}-z_{\min}.$
Here, $l_{x},l_{y},l_{z}$ represent the object size along each coordinate axis.
If we need the distance between two objects (or between the camera and an object), let $d_{12}$ be the Euclidean distance between their center points:
$d_{12}=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}+(z_{2}-z_{1})^{2}}.$ (2)
Here, $(x_{1},y_{1},z_{1})$ and $(x_{2},y_{2},z_{2})$ are the center coordinates of the two objects.
Spatial Relation
When the difference along one coordinate axis is significantly larger than along others, the sign of that difference determines the spatial relation:
$r_{xy}$ $=\mathrm{sign}(x_{A}-x_{B}),$ (3)
$r_{yz}$ $=\mathrm{sign}(y_{A}-y_{B}),$
$r_{zx}$ $=\mathrm{sign}(z_{A}-z_{B}).$
Here, $r_{xy},r_{yz},r_{zx}$ indicate relative positioning along each axis (e.g., front/back, left/right, above/below). We choose the axis with the greatest difference to label the dominant relation.
3D Video Grounding
For frame $t$ in the camera coordinate system, the 3D bounding box of an object can be described with dimensions, center position, and optional rotations:
$\text{BBox}_{t}=(\,l_{t},\,w_{t},\,h_{t},\,x_{t},\,y_{t},\,z_{t},\,\text{yaw}_{t},\,\text{pitch}_{t},\,\text{roll}_{t}\,).$ (4)
Here, $(l_{t},w_{t},h_{t})$ are the object dimensions, $(x_{t},y_{t},z_{t})$ is the center position, and $(\text{yaw}_{t},\text{pitch}_{t},\text{roll}_{t})$ are optional rotation angles if available.
A.2.2 Dynamic Understanding
Pose Estimation
Given the camera’s initial pose $(p_{0},o_{0})$, the pose $(p_{t},o_{t})$ at time $t$ can be obtained using the extrinsic-derived matrices $R_{t}$ (rotation) and $T_{t}$ (translation):
$p_{t}$ $=R_{t}\,p_{0}+T_{t},$ (5)
$o_{t}$ $=o_{0}+\Delta o_{t}.$
Here, $p_{t}$ is the position, $o_{t}$ is the orientation.
Displacement and Path Length
Let $p_{i}=(x_{i},y_{i},z_{i})$ be the position at time $i$. The displacement $d_{0n}$ and path length $L_{\text{traj}}$ are computed as:
$d_{0n}=\sqrt{(x_{n}-x_{0})^{2}+(y_{n}-y_{0})^{2}+(z_{n}-z_{0})^{2}},$ (6)
$L_{\text{traj}}=\sum_{i=1}^{n}\sqrt{(x_{i}-x_{i-1})^{2}+(y_{i}-y_{i-1})^{2}+(z_{i}-z_{i-1})^{2}}.$ (7)
Here, $d_{0n}$ is the straight-line distance from the initial to the final position; $L_{\text{traj}}$ sums consecutive segment lengths for the entire path.
Speed and Acceleration
Let $\Delta t$ be the time interval between consecutive frames. Then the speed $v_{i}$ and acceleration $a_{i}$ are:
$v_{i}$ $=\frac{d_{i}}{\Delta t},$ (8)
$a_{i}$ $=\frac{v_{i}-v_{i-1}}{\Delta t}.$
Here, $d_{i}$ is the displacement between adjacent frames, $v_{i}$ is the speed at time $i$, and $a_{i}$ is the acceleration.
Ego-Centric Orientation
If $\theta_{t}$ denotes the camera orientation (azimuth) at time $t$, then the orientation change $\Delta\theta_{t}$ is:
$\Delta\theta_{t}=\theta_{t}-\theta_{0}.$ (9)
This indicates how much the camera has rotated relative to its initial azimuth.
Trajectory Description
We apply the Ramer-Douglas-Peucker (RDP) algorithm to simplify the sequence of positions into key line segments. The resulting polyline is described in a piecewise manner (e.g., “go straight for 30m, turn left 85°, then go straight for 20m, …”), providing a concise representation of complex motion trajectories.
A.3 Human Involved Quality Control
Despite leveraging automated methods for generating large quantities of question-answer pairs, several critical issues necessitated human intervention for quality assurance. These included inaccuracies or ambiguities in automatically generated object descriptions, logical inconsistencies or irrelevant content within questions and options, and insufficient visual information in video data due to occlusions or limited resolution. To address these challenges, we established a dedicated online platform for systematic multi-round manual review and random sampling checks. Human reviewers evaluated each QA pair based on several criteria: clarity and accuracy of object descriptions, logical coherence between questions and answer options, sufficiency of visual information to accurately respond to questions, and effectiveness of distractors in posing genuine challenges. After passing the rigorous review process, answer options were randomized to prevent bias arising from fixed option ordering. Ultimately, this meticulous procedure enabled us to select over 2,000 high-quality QA pairs from more than 300 video sequences, significantly enhancing the robustness and reliability of the STI-Bench benchmark dataset.
Appendix B Evaluation Details
The evaluation framework of STI-Bench is carefully designed to assess multimodal models across diverse real-world scenarios, including Desktop, Indoor, and Outdoor environments. Specifically, the benchmark incorporates tasks derived from three publicly available datasets: Omni6DPose for desktop-scale, ScanNet for indoor, and Waymo for outdoor. Each task comprises question-answer pairs generated using Multimodal Large Language Models, which provide detailed semantic descriptions for objects based on video frames and corresponding multimodal data.
To ensure task diversity and challenge, each question is accompanied by carefully constructed distractors, differing significantly yet plausibly from the ground truth. Recognizing the variance in precision requirements among different scenarios—millimeter-level for Desktop, centimeter-to-decimeter-level for Indoor, and decimeter-to-meter-level for Outdoor—we apply scenario-specific scaling factors to generate distractors. Logarithmic sampling techniques are also utilized to ensure an even distribution of numerical differences, thus maintaining evaluation robustness.
Furthermore, rigorous human quality control was conducted, involving multiple rounds of manual filtering and random sampling reviews. This step ensured the logical consistency, clarity, and accuracy of all QA pairs. Finally, answer options were randomized post-review to prevent ordering bias. As a result, the final benchmark contains more than 2,000 meticulously curated QA pairs, providing a comprehensive and reliable evaluation framework across diverse real-world scenarios.
Appendix C Visualization Results
To qualitatively illustrate the effectiveness and interpretability of STI-Bench, visualization examples from representative Desktop, Indoor, and Outdoor scenarios are presented. Visualizations include video frames with annotated bounding boxes and corresponding object descriptions generated by Multimodal Large Language Models (MLLMs). For instance, objects are annotated with clear descriptions such as “a beige minivan with a roof rack” or “a red backpack on a brown leather sofa,” which are directly visible in the annotated frames. Additionally, visual representations of object trajectories, camera movements, and bounding box projections are provided, clearly demonstrating the spatial relationships and dynamic interactions within each scenario. These visualizations not only validate the accuracy and robustness of generated annotations but also provide intuitive insights into the model’s interpretability and performance in diverse real-world settings. Additionally, we also showcase typical examples of the Gemini 2.5 Pro’s thought process as used during the error analysis.
Figure 8: STI-Bench Examples (Outdoor)
Figure 9: STI-Bench Examples (Indoor)
Figure 10: STI-Bench Examples (Desktop)
Figure 11: Thinking example (Correct)
Figure 12: Thinking example (Incorrect)